OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Create an equation, then use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. Equation: Explanation: Solution:

OpenStudy (anonymous):

P = the profit they make n = the number of flowers they sell m = the mark-up of the flowers c = the cost of a flower Then P = mcn - cn Rearrange to find n P = n(mc - c) n = P / (mc - c) Substituting values n = 500 / (1.4*0.5 - 0.5) n = 2500 They must sell 2500 flowers to make a profit of $500.

OpenStudy (anonymous):

also so they sell a flower at 140% from the original price that would be 0.50$......100% x$.............140% x= 0.50*140/100=0.70$ a flower to make 500$ you should sell 500/0.70= 714 flowers

OpenStudy (anonymous):

i saw that one on someones openstudy question i jinda wanted a difrent answer

OpenStudy (anonymous):

its not really helping when you copy some ones answers lol i already look at the other one where you copied the answers from

OpenStudy (anonymous):

lol now she leaves hahaha